# Dependency-aware Maximum Likelihood Estimation for Active Learning

### v1

@@ Abstract
Active learning aims to efficiently build a labeled training set by strategically selecting samples to query labels from annotators. In this sequential process, each sample acquisition influences subsequent selections, causing dependencies among samples in the labeled set. However, these dependencies are overlooked during the model parameter estimation stage when updating the model using Maximum Likelihood Estimation (MLE), a conventional method that assumes independent and identically distributed (i.i.d.) data. We propose Dependency-aware MLE (DMLE), which corrects MLE within the active learning framework by addressing sample dependencies typically neglected due to the i.i.d. assumption, ensuring consistency with active learning principles in the model parameter estimation process. This improved method achieves superior performance across multiple benchmark datasets, reaching higher performance in earlier cycles compared to conventional MLE. Specifically, we observe average accuracy improvements of 6%, 8.6%, and 10.5% for $k=1$, $k=5$, and $k=10$ respectively, after collecting the first 100 samples, where entropy is the acquisition function and $k$ is the query batch size acquired at every active learning cycle.
@@ 1 Introduction
Active learning is an iterative process that alternates between two steps: (1) querying samples/labels and (2) adapting or retraining the model with the updated training set (Settles, 2009) as illustrated in Figure 2. Each complete iteration of these steps is referred to as a cycle. In each cycle, the model might either acquire one sample or a batch of samples during the query sampling step. So, in an active learning setup, previously selected data have control over the selection of new samples at each cycle through the updated model parameters. This sequential nature of data collection in active learning inherently introduces dependencies between samples across cycles in the labeled set. However, conventional Maximum Likelihood Estimation (MLE) assumes samples are i.i.d., overlooking dependencies during model parameter estimation; hence, we refer to it as Independent MLE (IMLE).
Sample dependency has been studied in the literature by revisiting the sample acquisition stage (Sener & Savarese, 2018; Ash et al., 2019a) or re-weighting the objective function (Farquhar et al., 2021; Beygelzimer et al., 2008), however, these approaches are insufficient to capture the dependency within the objective function, as they still rely on the i.i.d. assumption of MLE. Neglecting sample dependency during model parameter estimation is fundamentally incompatible with the nature of active learning. Ignoring this dependency during training leads to a vicious cycle of suboptimal model parameter estimation and, in turn, poor sample acquisition decisions in upcoming cycles as illustrated in Figure 1. The model updated with DMLE, using the same number of samples, outperforms the one updated with IMLE, creating a decision boundary similar to the one achieved by using all data samples. The diverse and well-separated sample selections made by the DMLE-updated model show that a better model leads to more effective sample selection and better results with fewer samples. This motivates us to explore ways to model this dependency in the training objective. To the best of our knowledge, this is the first work that attempts to correct for dependencies among samples across active learning cycles during model parameter estimation by removing the incompatible i.i.d. data assumption from the training objective. Our main contributions are as follows:
@@ 3 Background
We consider an active learning setup (Settles, 2009; Takezoe et al., 2023; Joshi et al., 2009), in which a large number of unlabeled samples are available for an evaluation in terms of an acquisition score and labeling by an oracle (e.g., an expert) on demand. An active learning algorithm proceeds in stages, collecting a batch of $k$ labels from the labeling oracle, training a model from the labeled data, and selecting which batch to collect next. At cycle $t$, we define the unlabeled dataset as $U_{t}=\{x_{i}\}_{i=1}^{N_{U}^{t}}$ and the labeled set as $D_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{L}^{t}}$ where $x_{i}\in\mathcal{X}\subset\mathbb{R}^{d_{x}}$ are the features, $y_{i}\in\mathcal{Y}=\{1,\ldots,K\}$ are categorical labels. We also denote by $N_{U}^{t}$ the number of unlabeled samples and $N_{L}^{t}$ the number of labeled samples at cycle $t$. We assume that, conditioned on parameters $\theta$ and sample features $x$, label $y$ is sampled from a conditional distribution $P(\cdot\mid x;\theta)$ parameterized by $\theta\in\mathbb{R}^{d_{\theta}}$. For example, in a deep neural-network (DNN) setting, where the DNN model is represented by function $f_{\theta}:\mathcal{X}\rightarrow\{0,1\}^{K}$, the probability would be given by the softmax function, i.e., $P(y\mid x,\theta)=e^{(f_{\theta}(x))_{y}}/\sum_{y^{\prime}}e^{(f_{\theta}(x))_{y^{\prime}}}$, for $y,y^{\prime}\in\mathcal{Y}$.
@@ 3.2 Sample Selection Strategy
where $\tilde{a}$ is a modified acquisition function and $\epsilon$ is Gumbel noise. This perturbation introduces probabilistic selection into the sample selection procedure. Kirsch et al. (2021) show that, given an acquisition function $a(\cdot)$, the sample selection distributions $P(x|D_{t};\theta)$ presented on Table 1 can be implemented via equation 6, through an appropriate choice of $\tilde{a}(\cdot)$ and parameters of the Gumbel noise. Details are provided in Appendix A in the supplement.
@@ 4 Dependency-aware Maximum Likelihood Estimation
However, this poses a challenge, as $P(S_{\tau}\mid D_{\tau-1};\theta)$ distribution is difficult to express in a closed form: it depends on the labels in $D_{t+1}$ through the generation of an estimate of $\theta$, which affects the acquisition function and in turn the sample selection. Fortunately, when utilizing the stochastic batch sampling approaches introduced by Kirsch et al. (2021) and discussed in Section 3.2, an appropriate corresponding probability distribution $P(S_{\tau}\mid D_{\tau-1};\theta)$ is associated with them (see Table 1). This allows us to introduce an appropriate correction term for equation 11. When using the Top-$k$ sampling discussed in Section 3.2, we approximate $P(S_{\tau}\mid D_{\tau-1};\theta)$ by the Softmax distribution given in Table 1 (Bishop, 2006).
@@
Let $P^{true}$ be the true data distribution, $(x^{\prime},y^{\prime})\sim P^{true}$, and $P(y^{\prime}\mid x^{\prime},D_{t};\hat{\theta})$ and $P(x^{\prime}\mid D_{t};\hat{\theta})$ be the distributions estimated from actively selected training data $D_{t}$. The expected log-likelihood on this data decomposes as:
$\mathbb{E}_{(x^{\prime},y^{\prime})\sim P^{true}}\Bigl{[}\ln L(\hat{\theta})\Bigr{]}=-D_{\text{KL}}(P(y^{\prime}\mid x^{\prime})\parallel P(y^{\prime}\mid x^{\prime},D_{t};\hat{\theta}))-D_{\text{KL}}(P(x^{\prime})\parallel P(x^{\prime}\mid D_{t};\hat{\theta})).$
When the training data is actively selected, DMLE, which includes both terms in the log-likelihood decomposition, minimizes the sum of two KL divergences, ensuring better alignment with the true distribution compared to IMLE, which ignores the dependency term. The dependency term, $D_{\text{KL}}\left(P(x^{\prime})\parallel P(x^{\prime}\mid D_{t};\hat{\theta})\right),$ measures the discrepancy between the true distribution $P(x^{\prime})$ and the model’s estimate $P(x^{\prime}\mid D_{t};\hat{\theta})$ learned from the actively selected data. Maximizing the expected log-likelihood is equivalent to minimizing the sum of two KL divergences, ensuring that both the predictive and data distributions of the model remain close to the true distributions. By addressing this discrepancy, DMLE reduces distributional mismatch and enhances the model’s approximation of the true data distribution. In contrast, IMLE, by neglecting this term, risks a greater distributional gap. The dependency term ensures that the model’s input distribution remains aligned with the true data distribution. The overall algorithm with DMLE in the active learning framework is provided in Alg. 1.
Algorithm 1 Maximum Likelihood Estimation for Active Learning
@@ 5.1 Experiment Setting
Datasets and Models. We evaluate the impact of including the dependency term during model parameter estimation for the active learning process on the model’s prediction performance by testing it on datasets consisting of images, texts, or features; to have diverse experiments, we used datasets with different sizes and complexities. We use eight classification datasets: Iris (Fisher, 1988), Mnist (Deng, 2012), Fashion-Mnist (FMnist)  (Xiao et al., 2017), SVHN (Netzer et al., 2011), Reuters-21578 (Reuters) (Lewis, 1997), Emnist-Letters (Emnist) (Cohen et al., 2017), Cifar-10 (Krizhevsky et al., 2009), and Tiny ImageNet (Le & Yang, 2015). We randomly sample a portion of each dataset and assume that only the subset was available during training. We used a subset of 30000 samples for Mnist and FMnist, 110 samples for Iris, 14651 samples for SVHN, 8083 samples for Reuters, 4440 samples for Emnist, 3000 samples for Cifar-10, and 1500 samples for Tiny ImageNet. Depending on the dataset, we adjust the complexity of the model. We used LeNet (Lecun et al., 1998) for Mnist, FMnist, SVHN, and Emnist, a two-layer MLP for Iris, a neural network with embedding and convolutional layers for Reuters, ResNet-50 (He et al., 2015) for Cifar-10, and a visual transformer (ViT-B/16) with 200 classes fine-tuned from ImageNet weights (Dosovitskiy et al., 2021).
@@
The complexity of model parameter estimation differs between IMLE and DMLE due to the dependency term in the DMLE objective function, which requires additional calculations per sample. We perform the complexity analysis per sample in an epoch, assuming the use of Stochastic Gradient Descent for gradient updates and disregarding its associated complexity. Under these conditions, the time complexity of IMLE is $\mathcal{O}(N_{L})$, while that of DMLE is $\mathcal{O}(cN_{L})$, where $N_{L}$ represents the number of labeled samples. Experimental timing results are provided in Appendix F.
@@ 5.2 Results
Table 2: Comparison of mean classification test accuracies ±$\pm 1$ standard deviation for different sample selection sizes ($k$) during cycles and different sample selection strategies where number of samples in the labeled set $N_{L}=100$. We use entropy for the acquisition function and SSMS (Stochastic Softmax Sampling), SPS (Stochastic Power Sampling), SSRS (Stochastic Soft-rank Sampling), and Top-$k$ sampling for the sample selection strategy. The bold text highlights the higher mean accuracy and lower standard deviation when comparing DMLE and IMLE under various sampling schemes.
@@
SVHN 1 $\mathbf{0.30\pm 0.02}$ $0.27\pm 0.03$ $\mathbf{0.30\pm 0.02}$ $0.28\pm 0.03$ $\mathbf{0.25\pm 0.03}$ $0.23\pm 0.03$ $\mathbf{0.22\pm 0.01}$ $0.17\pm 0.03$
5 $\mathbf{0.27\pm 0.05}$ $0.23\pm 0.02$ $\mathbf{0.25\pm 0.04}$ $0.24\pm 0.03$ $\mathbf{0.22\pm 0.03}$ $0.20\pm 0.04$ $\mathbf{0.22\pm 0.02}$ $0.19\pm 0.02$
@@
With the experiments, we seek to highlight the impact of not eliminating the dependency term during the model parameter estimation. Thus, the experiments present the test performance discrepancy between using DMLE vs. IMLE for multiple benchmark datasets with various sample selection schemes and acquisition functions. We provide the experiment results with the entropy and Core-set acquisition functions below. A similar case study for BALD and least confidence acquisition functions is presented in Appendix E.
Table 2 demonstrates the test accuracy results of using DMLE vs. IMLE after 100 collected samples across seven datasets, employing entropy as the acquisition function, with four different sample selection schemes—Stochastic Softmax Sampling (SSMS), Stochastic Power Sampling (SPS), Stochastic Soft-rank Sampling (SSRS), and Top-$k$ sampling—along with batch sizes $k\in\{1,5,10\}$. When SSMS or SPS are used for the sample selection, DMLE leads to better test accuracy results for all datasets and batch sizes except for the Iris dataset with $k=1$. In the case of SSRS as the sample selection strategy, test accuracies are either improved when DMLE is employed as the method of model parameter estimation or competitive with IMLE, except for FMnist with $k=1$ or $k=10$. For Top-$k$ selection, DMLE either outperforms or leads to comparable results with IMLE in all cases except Emnist with $k=5$, FMnist with $k=1$, and Cifar-10 with $k=1$. The average test accuracy improvements of 9.6%, 7%, 3.9%, and 11.8% are observed for SSMS, SPS, SSRS, and Top-$k$ sampling strategies respectively, which underline the significance of sample dependencies in model parameter estimation. Using different $k$ values for the sample selection size yields average test accuracy improvements of 6%, 8.6%, and 10.5% for $k=1$, $k=5$, and $k=10$ respectively. We observe performance degradation as the sample selection size $k$ increases for both DMLE and IMLE. We conjecture that this drop occurs due to the reduced diversity in the collected samples, where similar samples might have been collected in an active learning cycle. Even though we observe a trend for the test accuracy to drop when $k$ is increased, the reported average accuracy improvements for different $k$ values suggest that DMLE leads to better results than IMLE. The experiments in Appendix E) further underline the importance of using DMLE in active learning for model parameter estimation and its impact on model performance.
@@
Figure 3: The average test accuracy comparison with ±$\pm 1$ standard deviation for DMLE and IMLE over cycles for Iris, SVHN, Reuters, and Tiny ImageNet datasets for different sample selection strategies, namely Stochastic Softmax Sampling(SSMS), Stochastic Power Sampling(SPS), Stochastic Soft-rank Sampling(SSRS), and Top-$k$ Sampling where sample selection size $k=1$ for all except Tiny ImageNet and $k=5$ for Tiny ImageNet.
To analyze the accuracy evolution over the active learning cycles we plot the mean test accuracy ±$\pm 1$ standard deviation for DMLE and IMLE using four sample selection strategies with entropy acquisition function. Figure 3 shows such curves for the Iris, SVHN, Reuters, and Tiny ImageNet datasets where the sample selection size is $k=1$ for the first three and $k=5$ for Tiny ImageNet, with the remaining four datasets (MNIST, Fashion-Mnist, Emnist, and Cifar-10) provided in Appendix E. In the first row of Figure 3, the test accuracy performance for the Iris dataset after 100 cycles of collecting one sample at each cycle is shown for different sample selection schemes. The initial unlabeled set size for the Iris dataset is 110 which means after 100 cycles, we collected almost all of the samples in the unlabeled set. Thus, it is observed that DMLE and IMLE exhibit similar performance in the later cycles, while DMLE yields better test accuracy in the earlier cycles, demonstrating its advantage within the active learning framework. In the second and third rows of Figure 3, the average test accuracy plots for 500 cycles on SVHN and Reuters datasets are given. The improvement gained by accounting for sample dependencies during model parameter estimation with DMLE is evident in all plots, regardless of the sample selection strategy employed. In the last row of Figure 3, the performance difference between DMLE and IMLE is evident across all selection strategies for the Tiny ImageNet dataset, with DMLE showing a particularly greater improvement when Top-$k$ selection is used.

### v2

@@ Abstract
Active learning aims to efficiently build a labeled training set by strategically selecting samples to query labels from annotators. In this sequential process, each sample acquisition influences subsequent selections, causing dependencies among samples in the labeled set. However, these dependencies are overlooked during the model parameter estimation stage when updating the model using Maximum Likelihood Estimation (MLE), a conventional method that assumes independent and identically distributed (i.i.d.) data. We propose Dependency-aware MLE (DMLE), which corrects MLE within the active learning framework by addressing sample dependencies typically neglected due to the i.i.d. assumption, ensuring consistency with active learning principles in the model parameter estimation process. This improved method achieves superior performance across multiple benchmark datasets, reaching higher performance in earlier cycles compared to conventional MLE. Specifically, we observe average accuracy improvements of 6%, 8.6%, and 10.5% for $k=1$, $k=5$, and $k=10$ respectively, after collecting the first 100 samples, where entropy is the acquisition function and $k$ is the query batch size acquired at every active learning cycle. Our implementation is publicly available at: https://github.com/neu-spiral/DMLEforAL
@@ 1 Introduction
Active learning is an iterative process that alternates between two steps: (1) querying samples/labels and (2) adapting or retraining the model with the updated training set (Settles, 2009). Each complete iteration of these steps is referred to as a cycle. In each cycle, the model might either acquire one sample or a batch of samples during the query sampling step. So, in an active learning setup, previously selected data have control over the selection of new samples at each cycle through the updated model parameters. This sequential nature of data collection in active learning inherently introduces dependencies between samples across cycles in the labeled set. However, conventional Maximum Likelihood Estimation (MLE) assumes samples are i.i.d., overlooking dependencies during model parameter estimation; hence, we refer to it as Independent MLE (IMLE).
Sample dependency has been studied in the literature by revisiting the sample acquisition stage (Sener & Savarese, 2018; Ash et al., 2019a) or re-weighting the objective function (Farquhar et al., 2021; Beygelzimer et al., 2008), however, these approaches are insufficient to capture the dependency within the objective function, as they still rely on the i.i.d. assumption of MLE. Neglecting sample dependency during model parameter estimation is fundamentally incompatible with the nature of active learning. Ignoring this dependency during training leads to a vicious cycle of suboptimal model parameter estimation and, in turn, poor sample acquisition decisions in upcoming cycles as illustrated in Figure 1. The model updated with DMLE, using the same number of samples, outperforms the one updated with IMLE, creating a decision boundary similar to the one achieved by using all data samples with passive learning. The diverse and well-separated sample selections made by the DMLE-updated model show that a better model leads to more effective sample selection and better results with fewer samples. This motivates us to explore ways to model this dependency in the training objective. To the best of our knowledge, this is the first work that attempts to correct for dependencies among samples across active learning cycles during model parameter estimation by removing the incompatible i.i.d. data assumption from the training objective. Our main contributions are as follows:
@@ 3 Background
We consider an active learning setup (Settles, 2009; Takezoe et al., 2023; Joshi et al., 2009), in which a large number of unlabeled samples are available for an evaluation in terms of an acquisition score and labeling by an oracle (e.g., an expert) on demand. An active learning algorithm proceeds in stages, collecting a batch of $k$ labels from the labeling oracle, training a model from the labeled data, and selecting which batch to collect next. At cycle $t$, we define the unlabeled dataset as $U_{t}=\{x_{i}\}_{i=1}^{N_{U}^{t}}$ and the labeled set as $D_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{L}^{t}}$ where $x_{i}\in\mathcal{X}\subset\mathbb{R}^{d_{x}}$ are the features, $y_{i}\in\mathcal{Y}=\{1,\ldots,K\}$ are categorical labels. We also denote by $N_{U}^{t}$ the number of unlabeled samples and $N_{L}^{t}$ the number of labeled samples at cycle $t$. We assume that, conditioned on parameters $\theta$ and sample features $x$, label $y$ is sampled from a conditional distribution $P(\cdot\mid x;\theta)$ parameterized by $\theta\in\mathbb{R}^{d_{\theta}}$. For example, in a deep neural-network (DNN) setting, where the DNN model is represented by function $f_{\theta}:\mathcal{X}\rightarrow\{0,1\}^{K}$, the probability would be given by the softmax function, i.e., $P(y\mid x,\theta)=e^{(f_{\theta}(x))_{y}}/\sum_{y^{\prime}}e^{(f_{\theta}(x))_{y^{\prime}}}$, for $y,y^{\prime}\in\mathcal{Y}$. For clarity, a glossary of the terminology is included in Appendix J.
@@ 3.2 Sample Selection Strategy
[…]
@@ 4 Dependency-aware Maximum Likelihood Estimation
However, this poses a challenge, as $P(S_{\tau}\mid D_{\tau-1};\theta)$ distribution is difficult to express in a closed form: it depends on the labels in $D_{t+1}$ through the generation of an estimate of $\theta$, which affects the acquisition function and in turn the sample selection. Fortunately, when utilizing the stochastic batch sampling approaches introduced by Kirsch et al. (2021) and discussed in Section 3.2, an appropriate corresponding probability distribution $P(S_{\tau}\mid D_{\tau-1};\theta)$ is associated with them (see Table 1). This allows us to introduce an appropriate correction term for equation 11. When using the Top-$k$ sampling discussed in Section 3.2, we approximate $P(S_{\tau}\mid D_{\tau-1};\theta)$ by the Softmax distribution given in Table 1 (Bishop, 2006), with empirical support for this choice provided in Appendix I.
Unlike the formulations presented by Raginsky & Rakhlin (2011) (equation 11) and discussed in Chap. 15 of Lattimore & Szepesvári (2019), our approach explicitly models both the data likelihood and the sample selection distribution as functions of a single underlying parameter $\theta$. This choice reflects the interactive nature of active learning, where the evolving model directly influences which samples are selected, and the acquired labels in turn refine the model. By updating the selection distribution at each cycle using the most recent parameter estimates, our formulation ensures that past acquisitions are re-evaluated under the current model, capturing the feedback loop between selection and learning. This stands in contrast to approaches that treat the selection mechanism as fixed or independent of the model dynamics, which neglect the dependencies inherent to the sequential data generation process in active learning.
@@
[…]
$).\mathbb{E}_{(x^{\prime},y^{\prime})\sim P^{true}}\Bigl[\ln L(\hat{\theta})\Bigr]=-D_{\text{KL}}(P(y^{\prime}\mid x^{\prime})\parallel P(y^{\prime}\mid x^{\prime},D_{t};\hat{\theta}))-D_{\text{KL}}(P(x^{\prime})\parallel P(x^{\prime}\mid D_{t};\hat{\theta}$)).
[…]
Algorithm 1 Dependency-aware Maximum Likelihood Estimation for Active Learning
@@ 5.1 Experiment Setting
Datasets and Models. We evaluate the impact of including the dependency term during model parameter estimation for the active learning process on the model’s prediction performance by testing it on datasets consisting of images, texts, or features; to have diverse experiments, we used datasets with different sizes and complexities. We use eight classification datasets: Iris (Fisher, 1988), Mnist (Deng, 2012), Fashion-Mnist (FMnist)  (Xiao et al., 2017), SVHN (Netzer et al., 2011), Reuters-21578 (Reuters) (Lewis, 1997), Emnist-Letters (Emnist) (Cohen et al., 2017), Cifar-10 (Krizhevsky et al., 2009), and Tiny ImageNet (Le & Yang, 2015). We randomly sample a portion of each dataset and assume that only the subset was available during training. We used a subset of 30000 samples for Mnist and FMnist, 110 samples for Iris, 14651 samples for SVHN, 8083 samples for Reuters, 4440 samples for Emnist, 3000 samples for Cifar-10, and 1500 samples for Tiny ImageNet. Depending on the dataset, we adjust the complexity of the model. We used LeNet (Lecun et al., 1998) for Mnist, FMnist, SVHN, and Emnist, a two-layer MLP for Iris, a neural network with embedding and convolutional layers for Reuters, ResNet-50 (He et al., 2015) for Cifar-10, and a visual transformer (ViT-B/16) with 200 classes fine-tuned from ImageNet weights (Dosovitskiy et al., 2021) for Tiny ImageNet.
@@
The complexity of model parameter estimation differs between IMLE and DMLE due to the dependency term in the DMLE objective function, which requires additional calculations per sample. We perform the complexity analysis per sample in an epoch, assuming the use of Stochastic Gradient Descent for gradient updates and disregarding its associated complexity. Under these conditions, the time complexity of IMLE is $\mathcal{O}(N_{L})$, while that of DMLE is $\mathcal{O}(cN_{L})$, where $N_{L}$ represents the number of labeled samples. Experimental timing results are provided in Appendix E.
@@ 5.2 Results
Table 2: Comparison of mean classification test accuracies $\pm 1$ standard deviation for different sample selection sizes ($k$) during cycles and different sample selection strategies where number of samples in the labeled set $N_{L}=100$. We use entropy for the acquisition function and SSMS (Stochastic Softmax Sampling), SPS (Stochastic Power Sampling), SSRS (Stochastic Soft-rank Sampling), and Top-$k$ sampling for the sample selection strategy. The bold text highlights the higher mean accuracy and lower standard deviation when comparing DMLE and IMLE under various sampling schemes.
@@
[…]
@@
With the experiments, we seek to highlight the impact of not eliminating the dependency term during the model parameter estimation. Thus, the experiments present the test performance discrepancy between using DMLE vs. IMLE for multiple benchmark datasets with various sample selection schemes and acquisition functions. We provide the experiment results with the entropy and Core-set acquisition functions in this section. A similar case study for BALD and least confidence acquisition functions is presented in Appendix F.
Table 2 demonstrates the test accuracy results of using DMLE vs. IMLE after 100 collected samples across seven datasets, employing entropy as the acquisition function, with four different sample selection schemes—Stochastic Softmax Sampling (SSMS), Stochastic Power Sampling (SPS), Stochastic Soft-rank Sampling (SSRS), and Top-$k$ sampling—along with batch sizes $k\in\{1,5,10\}$. When SSMS or SPS are used for the sample selection, DMLE leads to better test accuracy results for all datasets and batch sizes except for the Iris dataset with $k=1$. In the case of SSRS as the sample selection strategy, test accuracies are either improved when DMLE is employed as the method of model parameter estimation or competitive with IMLE, except for FMnist with $k=1$ or $k=10$. For Top-$k$ selection, DMLE either outperforms or leads to comparable results with IMLE in all cases except Emnist with $k=5$, FMnist with $k=1$, and Cifar-10 with $k=1$. The average test accuracy improvements of 9.6%, 7%, 3.9%, and 11.8% are observed for SSMS, SPS, SSRS, and Top-$k$ sampling strategies respectively, which underline the significance of sample dependencies in model parameter estimation. Using different $k$ values for the sample selection size yields average test accuracy improvements of 6%, 8.6%, and 10.5% for $k=1$, $k=5$, and $k=10$ respectively. We observe performance degradation as the sample selection size $k$ increases for both DMLE and IMLE. We conjecture that this drop occurs due to the reduced diversity in the collected samples, where similar samples might have been collected in an active learning cycle. Even though we observe a trend for the test accuracy to drop when $k$ is increased, the reported average accuracy improvements for different $k$ values suggest that DMLE leads to better results than IMLE. The experiments in Appendix F) further underline the importance of using DMLE in active learning for model parameter estimation and its impact on model performance.
@@
Figure 3: The average test accuracy comparison with $\pm 1$ standard deviation for DMLE and IMLE over cycles for Iris, SVHN, Reuters, and Tiny ImageNet datasets for different sample selection strategies, namely Stochastic Softmax Sampling(SSMS), Stochastic Power Sampling(SPS), Stochastic Soft-rank Sampling(SSRS), and Top-$k$ Sampling where sample selection size $k=1$ for all except Tiny ImageNet and $k=5$ for Tiny ImageNet.
To analyze the accuracy evolution over the active learning cycles we plot the mean test accuracy $\pm 1$ standard deviation for DMLE and IMLE using four sample selection strategies with entropy acquisition function. Figure 3 shows such curves for the Iris, SVHN, Reuters, and Tiny ImageNet datasets where the sample selection size is $k=1$ for the first three and $k=5$ for Tiny ImageNet, with the remaining four datasets (MNIST, Fashion-Mnist, Emnist, and Cifar-10) provided in Appendix F. In the first row of Figure 3, the test accuracy performance for the Iris dataset after 100 cycles of collecting one sample at each cycle is shown for different sample selection schemes. The initial unlabeled set size for the Iris dataset is 110 which means after 100 cycles, we collected almost all of the samples in the unlabeled set. Thus, it is observed that DMLE and IMLE exhibit similar performance in the later cycles, while DMLE yields better test accuracy in the earlier cycles, demonstrating its advantage within the active learning framework. In the second and third rows of Figure 3, the average test accuracy plots for 500 cycles on SVHN and Reuters datasets are given. The improvement gained by accounting for sample dependencies during model parameter estimation with DMLE is evident in all plots, regardless of the sample selection strategy employed. In the last row of Figure 3, the performance difference between DMLE and IMLE is observable across all selection strategies for the Tiny ImageNet dataset, with DMLE showing a particularly greater improvement when Top-$k$ selection is used.